twenty seven drops of same sizes are charged at 220V each. They coalesce to form a bigger drop. Calculate the potential of the bigger drop.
1
Expert's answer
2020-04-15T10:10:02-0400

let r and R be radius of small and large drips respectively.

And v be the potential of small drip,

"v=220V"


Q and q be the charge on each small sphere and bigger sphere. and since charge is constant.


"Q=27q \to(1)"


From the constant volume of the drops,assuming drops are spherical applying volume equation,


"\frac{4}{3} \pi R^3=27\times\frac{4}{3} \pi r^3\\
R^3=(3r)^3\\
R=3r \to(2)"



Potential of bigger drop

"V_1=\frac{Q}{C}=\frac{27q}{3(4\pi\epsilon_ 0r)}=9[\frac{q}{4\pi \epsilon_0r}]\\
V_1=9\times220\\
V_1=1980V"